# Multi-Order Hyperbolic Graph Convolution and Aggregated Attention for Social Event Detection

### v1

@@ Abstract
Keywords Graph Neural Network, Hyperbolic Space, Social Event Detection
@@ 1 Introduction
Social Event Detection (SED) involves identifying interconnected message clusters from social media data streams to represent specific real-world events. One can view this process as a component of the broader social event data stream. Various areas, such as sentiment analysis, disaster response, voter trend prediction, and the detection of malicious events, widely apply SED. These applications are common across most mobile apps with social features, including major platforms like Twitter, Weibo, and Facebook. As a result, social event detection provides valuable insights not only for businesses, helping them assess consumer preferences, but also for public services, offering support in emergencies and disaster management. Given its broad utility, SED has garnered significant attention and discussion from both the academic and industrial sectors Fedoryszak et al. (2019); Peng et al. (2019); Zhou et al. (2020); Cao et al. (2021); Cui et al. (2021); Afyouni et al. (2022); Qian et al. (2023); Qiu et al. (2024).
Figure 1: (a) The points A and B in Euclidean space, situated on distinct branches of the diagram, are in greater proximity, but points A’ and B’ in hyperbolic space, also on different branches, exhibit a more reasonable closeness to one another. (b) A diagram illustrating the characteristics of hyperbolic space and its tangent plane, where lines e1 and e2 are two parallel lines passing through point c, which do not conform to the fifth axiom of contemporary geometry
SED presents unique challenges compared to traditional text classification tasks, primarily due to the nature of social media data. These challenges include short text length, unique data structures, and complex heterogeneous relationships.
The first challenge arises from the brevity of social media content. Time constraints, platform settings, or user preferences often constrain posts, resulting in short text that is challenging to process effectively. Creating high-quality labeled datasets from such content is also resource-intensive. Topic-based co-occurrence models were used in the early stages of SED research because they were effective and could show complex word relationships. However, these statistical methods suffered from sparse word representations and overlooked critical semantic information. To fix this problem, more recent research has used unsupervised or self-supervised methods like HISEvnet Cao et al. (2024)and HCRC Guo et al. (2024) to improve performance on high-quality datasets and added semantic context using tools like Word2Vec.
@@
  • Novel Framework: We propose MOHGCAA, a cutting-edge approach that integrates multi-order hyperbolic graph convolution with aggregated attention. This framework effectively addresses the limitations of traditional Euclidean-based methods by leveraging the hierarchical representation capabilities of hyperbolic space.
@@
  • Addressing Higher-Order Dependencies: MOHGCAA directly encodes higher-order syntactic relationships within hyperbolic space, overcoming the oversmoothing issue commonly associated with deep graph convolutional networks. This results in enriched feature diversity and improved performance, especially in scenarios involving complex, tree-like data structures.
@@ 2.1 Event Detection
Early event detection research focused on feature engineering to improve model performance. Ahn (2006) utilized constituent and dependency parsers to extract linguistic features like word forms, lemmas, and entity paths, supplemented by contextual information such as semantic roles. POS tagging provided grammatical cues for event triggers, while Liao and Grishman (2011) demonstrated the utility of unsupervised topic models in uncovering latent connections between events and entities, particularly in addressing imbalanced datasets like ACE-2005. Patwardhan and Riloff (2009) emphasized contextual features, using lexical and semantic patterns to enhance trigger-entity relationships. Despite their success, these approaches relied heavily on manual feature design, limiting scalability and paving the way for automated neural network methods. The advent of deep learning significantly advanced event detection. Nguyen and Grishman (2015) introduced convolutional neural networks (CNNs) for event detection, emphasizing feature extraction through pooling techniques. Chen et al. (2015) refined this with dynamic multi-pooling (DMCNN), segmenting sentences around event triggers to enhance semantic representation. Subsequent innovations, such as skipwindow CNNs (Zhang et al. (2016)) and parallel multi-pooling CNNs (Li et al. (2018)), improved compositional feature capture. Kodelja et al. (2019) extended CNNs to document-level contexts using a bootstrapping model. However, CNNs often struggled with long-range dependencies, as their localized focus and pooling operations sometimes led to information loss. To address these limitations, recurrent neural networks (RNNs), particularly Long Short-Term Memory (LSTM) networks (Hochreiter (1997)), became prominent. Ghaeini et al. (2016) and Chen et al. (2016) highlighted LSTMs’ ability to model sequential dependencies. Sha et al. integrated syntactic information using dependency bridges, blending hierarchical structures with sequential modeling. Li et al. (a) advanced this further with a Tree-LSTM framework enriched by external ontologies, enhancing long-range dependency modeling and event representation. Graph Convolutional Neural Networks (GCNs), introduced by Kipf and Welling (a), extended event detection beyond sequential representations by leveraging graph structures. GCN-ED (Nguyen and Grishman ) and JMEE (Liu et al. (a)) used dependency trees to model syntactic relationships. However, reliance on first-order syntactic edges limited these models’ ability to capture diverse relations. yan addressed this with the Multi-Order Graph Attention Network (MOGANED), enhancing representation diversity. Lai et al. proposed GatedGCN, introducing gating mechanisms to improve layer differentiation, while Cui et al. developed Edge-Enhanced GCN (EE-GCN) to incorporate dependency label information. Similarly, Dutta et al. improved event detection by integrating dependency edges in GTN-ED. The progress in social event detection builds upon the advancements in event detection research outlined above. By leveraging these foundational methodologies and adapting them to the unique challenges of social contexts, researchers have refined techniques to identify and analyze events within dynamic, interconnected social environments.
@@ 2.2 Social Event Detection
Social Event Detection (SED) has emerged as a challenging task that continues to garner substantial attention from researchers. Existing approaches to SED generally fall into two categories: topic-based methods and techniques leveraging heterogeneous information networks (HINs). Topic-Based Methods: Topic-based methods have been widely applied to social media data but face limitations due to the brevity and sparsity of text. For example, Latent Dirichlet Allocation (Blei et al. ) relies on word co-occurrence statistics, which are often insufficient in short-text scenarios. GPU-DMM Li et al. (b) mitigates this limitation by integrating Word2Vec with LDA and DMM, adding semantic background information to enhance performance. However, it overlooks relational structures between data nodes. To address this, SGNS Shi et al. models focus on learning contextual semantic relationships, effectively mitigating keyword sparsity in short texts. Inspired by SGNS, the Semantic-Assisted Non-Negative Matrix Factorization (SeaNMF) model was developed to capture semantics through word-document and word-context relevance in short-text corpora. While these methods have advanced topic modeling for SED, they generally fail to leverage the richness of HINs, which are capable of integrating multiple attribute representations to improve analytical depth. Heterogeneous Information Networks: The application of HINs to SED has shown considerable promise. Hao’s pioneering work introduced an event-based meta-pattern to capture semantic relevance among events, incorporating external knowledge bases to enrich event representation. Building on this, Hao proposed PP-GCN Peng et al. (2019), a heterogeneous graph convolutional network that integrates event metaschemas and external knowledge to achieve fine-grained event categorization. This model’s knowledge-instantiated similarity measure significantly enhanced its ability to identify semantic correlations among events, demonstrating exceptional performance in both detection and clustering tasks. Subsequent research has further exploited graph neural networks (GNNs) for SED. For instance, KPGNN, proposed by Cao et al. (a), utilizes incremental learning on HINs to improve detection accuracy. Despite these advances, existing methods struggle to accurately capture the hierarchical and tree-like structures inherent in social media data when constrained to Euclidean space. As highlighted earlier, this limitation poses a significant challenge for modeling multi-level relationships effectively. Our work addresses this gap by adopting hyperbolic space for node representation, enabling the modeling of multi-order relationships while preserving the hierarchical nature of social media data. This novel approach not only resolves structural representation issues but also offers a robust framework for advancing SED methodologies.
@@ 2.3 Hyperbolic Representation Learning
Recent years have witnessed a surge of research focused on modeling graphs in hyperbolic space. The primary motivation lies in the unique properties of hyperbolic space, which features negative curvature and geometrical advantages well-suited for learning hierarchical representations of graphs. Gromov demonstrated that hyperbolic space is inherently better than Euclidean space for representing tree-like structures. As noted earlier, GCN-based models improve syntactic and semantic modeling compared to conventional neural networks. Researchers Liu et al. (b) have extended graph neural networks to hyperbolic space by leveraging tangent space. For example, Peng et al. (a) constructed spatial-temporal graph convolutional networks in hyperbolic space for dynamic graph sequences, further exploring projection dimensions within tangent space using neural architecture search (NAS)Elsken et al. . Chami et al. (a) decoupled the message-passing process of GCN into three distinct operations: feature transformation, neighborhood aggregation, and activation. They then adapted each operation to hyperbolic space. Similarly, Bachmann et al. introduced k-GCN, a mathematically robust generalization of GCN for constant curvature spaces. In social event detection, hyperbolic space has been applied primarily to social event detection. Qiu et al. (a) demonstrated success using a simple HNN combined with MLP for social event detection. Building on this, Qiu et al. (b) proposed the GraphHAM model, which integrates hyperbolic space with an automatic meta-path selection mechanism for heterogeneous information graphs. This framework optimizes meta-path weights and converts them into vectors, significantly reducing the reliance on labeled data. Additionally, the authors designed a Hyperbolic Multi-Layer Perceptron (HMLP) to better capture semantic and structural information in social media contexts. Despite these advancements, current studies often overlook multi-order representations in hyperbolic space, leaving room for further innovation in event detection.
@@ 3.1 Definition of Graph Structure
A graph $G=(V,E,N,\mathcal{E},\mathcal{X})$ consists of nodes $V$, edges $E$, node subsets $N$, edge subsets $\mathcal{E}$, and node features $\mathcal{X}$. A graph is classified as heterogeneous if $|N|>1$ or $|\mathcal{E}|>1$, and homogeneous otherwise. Traditional Graph Neural Networks (GNNs) often use Euclidean space to model structural relationships, which can fall short in representing hierarchical information inherent to tree-like structures. Hyperbolic space, with its natural capacity for embedding hierarchical data, provides a more robust alternative, although it may still overlook complex inter-node relationships. This research addresses these limitations by adopting hyperbolic models, with the following section detailing relevant hyperbolic geometry concepts as a foundation.
@@ 3.2 Hyperbolic Models
Definition 3.1 (Poincaré Ball Model): The Poincaré Ball Model $\mathbb{P}^{n}_{c}$, characterized by negative curvature, is defined as the Riemannian manifold $(\mathcal{B}^{n}_{c},g^{\mathcal{B}}_{\mathbf{x}})$, where $\mathcal{B}^{n}_{c}=\{\mathbf{x}\in\mathbb{R}^{n}:\|\mathbf{x}\|^{2}<-1/c\}$ represents an open n-dimensional ball with radius $1/\sqrt{|c|}$. The metric tensor is given by $g^{B}_{\mathbf{x}}=(\lambda^{c}_{\mathbf{x}})^{2}g^{E}$, with $\lambda^{c}_{\mathbf{x}}=2/(1+c\|\mathbf{x}\|^{2})$ as the conformal factor, and $g^{E}$ as the Euclidean metric.
@@
This section investigates node representations for graph neural networks within hyperbolic space. In Euclidean space, node features are generally derived from pre-trained embeddings, node-specific attributes, random sampling, or one-hot encodings. In hyperbolic space, however, these features are projected into either the Poincaré ball model or the Lorentz model to accommodate the curvature of the space. In the Poincaré ball model, the node representation $\mathbf{x}^{\mathcal{B}}=\exp_{\mathbf{o}}^{c}(\mathbf{x}^{E})$ involves projecting features onto the tangent space at the origin. This projection is defined by:
@@
where $c$ denotes the curvature, $\mathbf{x}$ is a point in hyperbolic space, and $\mathbf{v}$ is a Euclidean vector. In the Lorentz model, the mapping is defined as: $\mathbf{x}^{\mathcal{L}}=\exp^{c}_{o}((0,\mathbf{x}^{E}))$, where a zero is prepended to $\mathbf{x}^{E}$ to satisfy the condition imposed by the Minkowski inner product. The exponential mapping in the Lorentz model is expressed as:
@@
After obtaining the initial hyperbolic space representation, it becomes necessary to perform feature transformation for graph neural networks. The most straightforward method is to complete Euclidean feature operations by projecting the existing hyperbolic space onto the tangent space at a specific point. The notations $\log_{x}^{c}(\mathbf{x}^{B})$ and $\log_{x}^{c}(\mathbf{x}^{C})_{[1:n]}$ represent operations in the Poincaré ball model and the Lorentz model, respectively. In the case of the Poincaré ball model, the mapping to the tangent space is defined as:
$\log_{\mathbf{x}}^{c}(\mathbf{y})=\frac{2}{\sqrt{|c|\lambda_{\mathbf{x}}^{c}}}\tanh^{-1}\left(\sqrt{|c|}\left\|-\mathbf{x}\oplus_{c}\mathbf{y}\right\|_{2}\right)\frac{-\mathbf{x}\oplus_{c}\mathbf{y}}{\left\|-\mathbf{x}\oplus_{\oplus_{c}}\mathbf{y}\right\|_{2}},$ (3)
@@
$\log_{\mathbf{x}}^{c}(\mathbf{y})=\frac{\cosh^{-1}(c\langle\mathbf{x},\mathbf{y}\rangle c)}{\sinh(\cosh^{-1}(c\langle\mathbf{x},\mathbf{y}\rangle))}(\mathbf{y}-c\langle\mathbf{x},\mathbf{y}\rangle_{\mathcal{L}}\mathbf{{x}}).$ (4)
After these operations, the resulting features undergo message aggregation and activation. They are then projected back to a new hyperbolic space using the exponential map, yielding a final hyperbolic space-based representation of the graph following a sequence of transformations. For further illustration, refer to Figure 4.

### v2

# Methodology, Writing – original draft, Writing – review & editing, Experiment

@@ Abstract
keywords:
[…]
{graphicalabstract}[Uncaptioned image]
{highlights}
We proposed MOHGCAA, a cutting-edge approach that integrates multi-order hyperbolic graph convolution with aggregated attention. This framework effectively addresses the limitations of traditional Euclidean-based methods by leveraging the hierarchical representation capabilities of hyperbolic space
The proposed framework undergoes rigorous testing across both supervised and unsupervised learning paradigms, demonstrating its adaptability and effectiveness across a range of tasks within hyperbolic space.
Extensive evaluations on a variety of benchmark datasets validate the robustness, scalability, and high performance of MOHGCAA. Its ability to capture multi-order relationships is particularly advantageous for hierarchical and tree-structured data.
MOHGCAA directly encodes higher-order syntactic relationships within hyperbolic space, overcoming the over-smoothing issue commonly associated with deep graph convolution networks. This results in enriched feature diversity and improved performance, especially in scenarios involving complex, tree-like data structures.
@@ 1 Introduction
Social Event Detection (SED) entails the identification of interconnected message clusters from social media data streams to signify certain real-world events. This procedure can be regarded as part of the larger social event data stream. SED is extensively utilized in diverse domains, including sentiment analysis, disaster response, voter trend prediction, and the identification of harmful occurrences. These applications are prevalent in the majority of mobile apps featuring social functionalities, encompassing prominent platforms such as Twitter, Weibo, and Facebook. Consequently, social event detection yields significant insights for both enterprises, aiding in the evaluation of consumer preferences, and public services, facilitating assistance during emergencies and disaster management. Due to its extensive applicability, SED has attracted considerable interest and discourse from both academic and industrial domains.
Figure 1: (a) The points A and B in Euclidean space, situated on distinct branches of the diagram, are in greater proximity, but points A’ and B’ in hyperbolic space, also on different branches, exhibit a more reasonable closeness to one another. (b) A diagram illustrating the characteristics of hyperbolic space and its tangent plane, where lines e1 and e2 are two parallel lines passing through point c, which do not conform to the fifth axiom of contemporary geometry.
SED presents unique challenges compared to traditional text classification tasks, primarily due to the nature of social media data. These challenges include short text length, unique data structures, and complex heterogeneous relationships [1, 2, 3, 4, 5, 6, 7, 8].
The first challenge arises from the brevity of social media content. Time constraints, platform settings, or user preferences often constrain posts, resulting in short text that is challenging to process effectively. Creating high-quality labeled datasets from such content is also resource-intensive. Topic-based co-occurrence models were used in the early stages of SED research because they were effective and could show complex word relationships. However, these statistical methods suffered from sparse word representations and overlooked critical semantic information. To fix this problem, more recent research has used unsupervised or self-supervised methods like HISEvnet [9] and HCRC [10] to improve performance on high-quality datasets and added semantic context using tools like Word2Vec [11].
@@
  1. 1.
    Novel Framework: We proposed MOHGCAA, a cutting-edge approach that integrates multi-order hyperbolic graph convolution with aggregated attention. This framework effectively addresses the limitations of traditional Euclidean-based methods by leveraging the hierarchical representation capabilities of hyperbolic space.
  2. 2.
@@
  4. 4.
    Addressing Higher-Order Dependencies: MOHGCAA directly encodes higher-order syntactic relationships within hyperbolic space, overcoming the over-smoothing issue commonly associated with deep graph convolution networks. This results in enriched feature diversity and improved performance, especially in scenarios involving complex, tree-like data structures.
@@ 2.1 Event Detection
Early event detection research focused on feature engineering to improve model performance. [12] utilized constituent and dependency parsers to extract linguistic features like word forms, lemmas, and entity paths, supplemented by contextual information such as semantic roles. POS tagging provided grammatical cues for event triggers, while [13] demonstrated the utility of unsupervised topic models in uncovering latent connections between events and entities, particularly in addressing imbalanced datasets like ACE-2005. [14] emphasized contextual features, using lexical and semantic patterns to enhance trigger-entity relationships. Despite their success, these approaches relied heavily on manual feature design, limiting scalability and paving the way for automated neural network methods. The advent of deep learning significantly advanced event detection. [15] introduced convolutional neural networks (CNNs) for event detection, emphasizing feature extraction through pooling techniques. [16] refined this with dynamic multi-pooling (DMCNN), segmenting sentences around event triggers to enhance semantic representation. Subsequent innovations, such as skip-window CNNs ([17]) and parallel multi-pooling CNNs ([18]), improved compositional feature capture. [19] extended CNNs to document-level contexts using a bootstrapping model. However, CNNs often struggled with long-range dependencies, as their localized focus and pooling operations sometimes led to information loss. To address these limitations, recurrent neural networks (RNNs), particularly Long Short-Term Memory (LSTM) networks ([20]), became prominent. [21] and [22] highlighted LSTMs’ ability to model sequential dependencies. [23] integrated syntactic information using dependency bridges, blending hierarchical structures with sequential modeling. [24] advanced this further with a Tree-LSTM framework enriched by external ontologies, enhancing long-range dependency modeling and event representation. Graph Convolution Neural Networks (GCNs), introduced by [25], extended event detection beyond sequential representations by leveraging graph structures. GCN-ED ([26]) and JMEE ([27]) used dependency trees to model syntactic relationships. However, reliance on first-order syntactic edges limited these models’ ability to capture diverse relations. [28] addressed this with the Multi-Order Graph Attention Network (MOGANED), enhancing representation diversity. [29] proposed GatedGCN, introducing gating mechanisms to improve layer differentiation, while [30] developed Edge-Enhanced GCN (EE-GCN) to incorporate dependency label information. Similarly, [31] improved event detection by integrating dependency edges in GTN-ED. The progress in social event detection builds upon the advancements in event detection research outlined above. By leveraging these foundational methodologies and adapting them to the unique challenges of social contexts, researchers have refined techniques to identify and analyze events within dynamic, interconnected social environments.
@@ 2.2 Social Event Detection
Social Event Detection (SED) has emerged as a challenging task that continues to garner substantial attention from researchers. Existing approaches to SED generally fall into two categories: topic-based methods and techniques leveraging heterogeneous information networks (HINs). Topic-Based Methods: Topic-based methods have been widely applied to social media data but face limitations due to the brevity and sparsity of text. For example, Latent Dirichlet Allocation [32] relies on word co-occurrence statistics, which are often insufficient in short-text scenarios. GPU-DMM [33] mitigates this limitation by integrating Word2Vec with LDA and DMM, adding semantic background information to enhance performance. However, it overlooks relational structures between data nodes. To address this, SGNS [34] models focus on learning contextual semantic relationships, effectively mitigating keyword sparsity in short texts. Inspired by SGNS, the Semantic-Assisted Non-Negative Matrix Factorization (SeaNMF) model was developed to capture semantics through word-document and word-context relevance in short-text corpora. While these methods have advanced topic modeling for SED, they generally fail to leverage the richness of HINs, which are capable of integrating multiple attribute representations to improve analytical depth. Heterogeneous Information Networks: The application of HINs to SED has shown considerable promise. Hao’s pioneering work introduced an event-based meta-pattern to capture semantic relevance among events, incorporating external knowledge bases to enrich event representation. Building on this, Hao proposed PP-GCN [2], a heterogeneous graph convolution network that integrates event metaschemas and external knowledge to achieve fine-grained event categorization. This model’s knowledge-instantiated similarity measure significantly enhanced its ability to identify semantic correlations among events, demonstrating exceptional performance in both detection and clustering tasks. Subsequent research has further exploited graph neural networks (GNNs) for SED. For instance, KPGNN, proposed by [4], utilizes incremental learning on HINs to improve detection accuracy. Despite these advances, existing methods struggle to accurately capture the hierarchical and tree-like structures inherent in social media data when constrained to Euclidean space. As highlighted earlier, this limitation poses a significant challenge for modeling multi-level relationships effectively. Our work addresses this gap by adopting hyperbolic space for node representation, enabling the modeling of multi-order relationships while preserving the hierarchical nature of social media data. This novel approach not only resolves structural representation issues but also offers a robust framework for advancing SED methodologies.
@@ 2.3 Hyperbolic Representation Learning
Recent years have witnessed a surge of research focused on modeling graphs in hyperbolic space. The primary motivation lies in the unique properties of hyperbolic space, which features negative curvature and geometrical advantages well-suited for learning hierarchical representations of graphs. Gromov demonstrated that hyperbolic space is inherently better than Euclidean space for representing tree-like structures. As noted earlier, GCN-based models improve syntactic and semantic modeling compared to conventional neural networks. Researchers [35] have extended graph neural networks to hyperbolic space by leveraging tangent space. For example, [36] constructed spatial-temporal graph convolution networks in hyperbolic space for dynamic graph sequences, further exploring projection dimensions within tangent space using neural architecture search (NAS). [37] decoupled the message-passing process of GCN into three distinct operations: feature transformation, neighborhood aggregation, and activation. They then adapted each operation to hyperbolic space. Similarly, [38] introduced k-GCN, a mathematically robust generalization of GCN for constant curvature spaces. In social event detection, hyperbolic space has been applied primarily to social event detection. [8] demonstrated success using a simple HNN combined with MLP for social event detection. Building on this, [39] proposed the GraphHAM model, which integrates hyperbolic space with an automatic meta-path selection mechanism for heterogeneous information graphs. This framework optimizes meta-path weights and converts them into vectors, significantly reducing the reliance on labeled data. Additionally, the authors designed a Hyperbolic Multi-Layer Perceptron (HMLP) to better capture semantic and structural information in social media contexts. Despite these advancements, current studies often overlook multi-order representations in hyperbolic space, leaving room for further innovation in event detection.
@@ 3 Preliminaries
Table 1: The notation used in this work.
Symbol Description
Space Symbol:
$\mathbb{E}$ Euclidean space.
$\mathbb{L}$ Hyperbolic space is represented by the Lorentz model.
$\mathbb{K}$ Hyperbolic space is represented by the Kelin model.
$\mathbb{B}$ Hyperbolic space is represented by the Poincaré Ball model.
Scalaers:
$\mathrm{c}$ The curvature of hyperbolic space ($\mathrm{c}$ < 0).
$1/\sqrt{|c|}$ The radius of an open n-dimensional ball.
$\lambda$ The conformal scaling factor.
vectors:
$\mathbf{v},\mathbf{x},\mathbf{y}$ Euclidean vectors.
$\mathbf{x}^{\mathcal{B}}$ The node representation in the Poincaré ball model.
$\mathbf{x}^{\mathcal{L}}$ The node representation in the Lorentz model.
$\mathbf{x}^{E}$ The Euclidean feature vector before hyperbolic mapping.
Matrices:
$A_{\mathrm{along}}$ The adjacency matrices with directed edges.
$A_{\mathrm{rev}}$ The adjacency matrices with the reverse direction of $A_{\mathrm{along}}$.
$A_{\mathrm{loop}}$ The adjacency matrices of an identity matrix for self-loops.
Operations:
[…]
${\langle\cdot,\cdot\rangle}_{L}$ The Minkowski inner product.
$\mathrm{diag}([...])_{n}$ The Minkowski metric.
$\exp_{\mathbf{o}}^{c}(\cdot)$ The exponential map on the hyperbolic manifold $\mathbb{H}_{c}^{d}$ with curvature $c$, computed at the origin $\mathbf{o}$.
$\log_{\mathbf{o}}^{c}(\cdot)$ The logarithmic map in the hyperbolic manifold $\mathbb{H}_{c}^{d}$ with curvature is $c$ computed at the origin $\mathbf{o}$.
$\oplus_{c}$ The Möbius addition (or hyperbolic addition) with the curvature of the hyperbolic space (often $c$ < 0)
$\oplus$ The element-wise addition.
$\kappa_{i,j}$ A learnable score that quantifies the importance of a neighbor $j$ to a node $i$ in the neighborhood aggregation.
$\otimes_{c}^{\mathcal{H}}$ The Möbius (hyperbolic) matrix multiplication operation with the curvature $c$.
$P_{o\to\mathbf{x}^{\mathcal{H}}}^{c}$ The parallel transport function from the origin $o$ to the hyperbolic point $\mathbf{x}^{\mathcal{H}}$ under curvature $c$.
@@ 3.1 Definition of Graph Structure
A graph $\mathcal{G}=(V,E,\mathcal{N},\mathcal{E})$ consists of node sets $V$, edge sets $E$, node type sets $\mathcal{N}$, and edge type sets $\mathcal{E}$. A graph is classified as heterogeneous if $\left|\mathcal{N}\right|>1$ or $\left|\mathcal{E}\right|>1$, and homogeneous otherwise. Traditional Graph Neural Networks (GNNs) often use Euclidean space to model structural relationships, which can fall short in representing hierarchical information inherent to tree-like structures. Hyperbolic space, with its natural capacity for embedding hierarchical data, provides a more robust alternative, although it may still overlook complex inter-node relationships. This research addresses these limitations by adopting hyperbolic models, with the following section detailing relevant hyperbolic geometry concepts as a foundation.
@@ 3.2 Hyperbolic Models
Definition 3.1 (Poincaré Ball Model): The Poincaré Ball Model $\mathbb{P}^{n}_{c}$, characterized by negative curvature, is defined as the Riemannian manifold $(\mathcal{B}^{n}_{c},g^{\mathcal{B}}_{\mathbf{x}})$, which $\mathcal{B}^{n}_{c}=\{\mathbf{x}\in\mathbb{R}^{n}:\|\mathbf{x}\|^{2}<-1/c\}$ represents an open n-dimensional ball with radius $1/\sqrt{|c|}$. The metric tensor is given by $g^{B}_{\mathbf{x}}=(\lambda^{c}_{\mathbf{x}})^{2}g^{E}$, with $\lambda^{c}_{\mathbf{x}}=2/(1+c\|\mathbf{x}\|^{2})$ as the conformal factor and $g^{E}$ as the Euclidean metric.
@@
This section investigates node representations for graph neural networks within hyperbolic space.
In Euclidean space, node features are generally derived from pre-trained embeddings, node-specific attributes, random sampling, or one-hot encodings. In hyperbolic space, however, these features are projected into either the Poincaré ball model or the Lorentz model to accommodate the curvature of the space.
In the Poincaré ball model, the node representation $\mathbf{x}^{\mathcal{B}}=\exp_{\mathbf{o}}^{c}(\mathbf{x}^{E})$ involves projecting features onto the tangent space at the origin. This projection is defined by:
@@
where $c$ denotes the curvature, $\mathbf{x}$ is a point in hyperbolic space, and $\mathbf{v}$ is a Euclidean vector.
In the Lorentz model, the mapping is defined as: $\mathbf{x}^{\mathcal{L}}=\exp^{c}_{o}((0,\mathbf{x}^{E}))$, where a zero is prepended $\mathbf{x}^{E}$ to satisfy the condition imposed by the Minkowski inner product. The exponential mapping in the Lorentz model is expressed as:
@@
After obtaining the initial hyperbolic space representation, it becomes necessary to perform feature transformation for graph neural networks. The most straightforward method is to complete Euclidean feature operations by projecting the existing hyperbolic space onto the tangent space at a specific point. The notations $\log_{x}^{c}(\mathbf{x}^{\mathcal{B}})$ and $\log_{x}^{c}(\mathbf{x}^{\mathcal{L}}))_{[1:n]}$ representation operations in the Poincaré ball model and the Lorentz model, respectively.
In the case of the Poincaré ball model, the mapping to the tangent space is defined as:
$\log_{\mathbf{x}}^{c}(\mathbf{y})=\frac{2}{\sqrt{|c|\lambda_{\mathbf{x}}^{c}}}\tanh^{-1}\left(\sqrt{|c|}\left\|-\mathbf{x}\oplus_{c}\mathbf{y}\right\|_{2}\right)\frac{-\mathbf{x}\oplus_{c}\mathbf{y}}{\left\|-\mathbf{x}\oplus_{c}\mathbf{y}\right\|_{2}}$ (3)
@@
$\log_{\mathbf{x}}^{c}(\mathbf{y})=\frac{\cosh^{-1}(c\langle\mathbf{x},\mathbf{y}\rangle c)}{\sinh(\cosh^{-1}(c\langle\mathbf{x},\mathbf{y}\rangle))}(\mathbf{y}-c\langle\mathbf{x},\mathbf{y}\rangle_{\mathcal{L}}\mathbf{{x}})$ (4)
. Figure 3 illustrates the mapping procedures between hyperbolic space and the corresponding Euclidean space.After these operations, the resulting features undergo message aggregation and activation. They are then projected back to a new hyperbolic space using the exponential map, yielding a final hyperbolic space-based representation of the graph following a sequence of transformations. For a more comprehensive illustration of the entire process described in this section, please refer to Figure 4.